### 17 Apr 2026 v1

@@ Abstract
We study a new measure of codependency in the second moment of a continuous-time multivariate asset price process, which we name the realized copula of volatility. The statistic is based on local volatility estimates constructed from high-frequency asset returns and affords a nonparametric estimator of the empirical copula of the latent stochastic volatility. We show consistency of our estimator with in-fill asymptotic theory, either with a fixed or increasing time span. In the latter setting, we derive a functional central limit theorem for the empirical process associated with the measurement error of the time-invariant marginal copula of volatility. We also develop a goodness-of-fit test to evaluate hypothesis about the shape of the latter. In a simulation study, we demonstrate that our estimator is a good proxy of both the empirical and marginal copula of volatility, even with a moderate amount of high-frequency data recorded over a relatively short sample. The goodness-of-fit test is found to exhibit size control and excellent power. We implement our framework on high-frequency transaction data from futures contracts that track the U.S. equity and treasury bond market. A Gumbel copula is found to offer a near-perfect bind between the realized variance processes in these data.
@@ 1 Introduction
Modeling the comovement of stochastic volatility across several assets is paramount to risk measurement and management, portfolio allocation, and multi-asset and derivative pricing. It has long been recognized that financial markets are interconnected, causing asset return volatilities—and correlations—to surge in tandem during financial distress, triggering significant volatility spillover effects, posing a potential systemic risk (see, e.g., Diebold and Yilmaz, 2009; Billio et al., 2012, and references therein). An understanding of volatility codependency is also important for pricing of volatility risk (i.e., the variance risk premium) across asset classes (Carr and Wu, 2009; Bollerslev and Todorov, 2011). The importance of the second moment structure of asset returns is further fueled by the so-called leverage effect that captures a negative association between return and volatility (Black, 1976; Christie, 1982). Thus, from a risk aggregation view the relevant target is often the dependence structure in second moment of returns (namely volatilities and correlations), rather than in the raw returns.
@@
In practice, the assessment of volatility comovements is complicated by the fact that volatility is unobserved, so it requires an empirical route based on an observable proxy. It therefore remains a relatively unexplored area. In recent years, however, the availability of high-frequency data has made it possible to estimate volatility nonparametrically with the realized variance (e.g., Andersen and Bollerslev, 1998; Barndorff-Nielsen and Shephard, 2002). Most of this literature concentrates on estimation of integrated measures of variation, defined as an integral of a smooth function of the latent spot variance process (see Jacod and Rosenbaum, 2013). However, as Li et al. (2013) point out, the mapping from the probability distribution of spot volatility to the distribution of integrated volatility is not one-to-one. Thus, the “integrated” approach may be partially informative about some distributional aspects of volatility and codependency, but it faces a limitation due to the inherent “identification failure.” This may weaken, or outright eliminate, power for testing against relevant alternatives. They suggest to work directly with the pathwise distribution of spot volatility via a so-called occupation measure (e.g., Geman and Horowitz, 1980), which recovers the empirical distribution function of the point-in-time volatility process (see also Li et al., 2016a, b). [^1] In addition, Christensen et al. (2019) and Zhang et al. (2022) made the univariate framework robust to microstructure noise. Regardless, the volatility occupation literature is entirely univariate and has concentrated on the marginal distribution, rather than looking at the interrelationship between various volatility processes.
@@
The starting point of our analysis is the remarkable result from probability theory known as Sklar’s theorem (Sklar, 1959). It states that for any continuous random vector $(X_{1},\dots,X_{d})$ with cumulative distribution function $H(x_{1},\dots,x_{d})=\mathbb{P}(X_{1}\leq x_{1},\dots,X_{d}\leq x_{d})$, there exists a unique function $C:[0,1]^{d}\rightarrow[0,1]$ called a copula, such that $H(x_{1},\dots,x_{d})=C(H_{1}(x_{1}),\dots,H_{d}(x_{d}))$, where $H_{i}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ is the marginal distribution of $X_{i}$. [^2] The concise mathematical definition is as follows: A $d$-dimensional copula $C:[0,1]^{d}\rightarrow[0,1]$ is the cumulative distribution function of a random vector $(U_{1},\dots,U_{d})$, where the marginal distribution of $U_{i}$ is uniform, i.e. $C(u_{1},\dots,u_{d})=\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ with $\mathbb{P}(U_{i}\leq u_{i})=u_{i}$, for $i=1,\dots,d$ and $0\leq u_{i}\leq 1$. Conversely, given $C(u_{1},\dots,u_{d})$ and $H_{1}(x_{1}),\dots,H_{d}(x_{d})$, $H(x_{1},\dots,x_{d})=C(H_{1}(x_{1}),\dots,H_{d}(x_{d}))$ is a probability distribution. Copulas provide a marginal-free representation of dependence in a multivariate setting, and they enable the researcher to model the marginal features of the data separately. [^3] The term copula originates from the Latin word for “bond” or “tie”. Indeed, many measures of concordance (and discordance), such as Kendall’s tau and Spearman’s rho, can be expressed entirely in terms of the copula (see Section 6). The latter are often more informative than the Pearson’s coefficient of correlation, which captures linear dependence. As emphasized by Embrects et al. (1999), correlation is insufficient for risk management. This is explained by the fact that financial markets typically exhibit nonlinear dependence and, further, may possess a high degree of tail concentration (or corner clustering). [^4] The application of copula theory had a revival in financial economics after Li (2000) proposed to model the default correlation in portfolios of bonds loans with a Gaussian copula. This enabled tractable analysis of structured credit products (such as collateralized debt obligations). The model became a standard pricing tool on Wall Street. However, since the Gaussian copula has no tail dependence, it tends to understate the probability of concurrent defaults, as seen during the financial crisis (see, for instance, Salmon, 2009, for a critique of the approach). A comprehensive review of the application of copula theory to economic and financial time series is Patton (2012). Therefore, copulas—or descriptive statistics derived from them—are a better representation of dependence. Another advantage of rank-based statistics is their invariance to monotonic transformations (that do not alter the ordering of the data). However, a model-free technique for analyzing the multivariate dependence structure of volatility is still lacking. In particular, tools for estimating and conducting inference on the copula of volatility are not available.
@@
We close that gap and propose to study the dependence structure of several stochastic volatility processes. We make a four-fold contribution to the literature. First, we extend Li et al. (2013) to the multivariate setting. We operate within a general, albeit standard, continuous-time arbitrage-free Itô semimartingale framework for a (multi-dimensional) log-price process, where the random volatility can exhibit rather unrestricted dynamics. We replace the latent spot variance processes by localized, jump-robust, realized variance measures constructed from discretely observed high-frequency returns over short time intervals. The estimation errors are controllable by the sampling interval and bandwidth selection (e.g., Jacod and Protter, 2012). We allow for very active jumps in the price process. The latter can be regarded as “nuisance parameters” in our setting, and we handle them with a truncation device (e.g., Mancini, 2009). Second, we construct a nonparametric estimator of volatility codependency, called the realized copula of volatility, enabling a meaningful cross-asset comparison irrespective of the marginal distribution of the individual volatility processes. Third, we develop the necessary large-sample theory for our estimator aligned with its target at each horizon. In the in-fill asymptotic limit with fixed time span, we show the realized copula of volatility converges, uniformly, in probability to its empirical counterpart that captures the volatility dependency over the time interval observed so far. Importantly, we allow for nonstationary volatility. In a double asymptotic in-fill and long-span setting—with stationary and weakly dependent volatility—we establish a functional central limit theorem in the form of weak convergence of the measurement error of the realized volatility copula. We show how this can be exploited to construct pointwise or, with a bit more effort, uniform confidence bands. A goodness-of-fit test for assessing the shape of the volatility copula is also proposed. The latter procedure can, for instance, be used to gauge the adequacy of various parametric copulas; an approach we look further at in our application. This builds upon and extends earlier results on the empirical copula process (Deheuvels, 1979; Fermanian et al., 2004). However, as explained above, our framework is distinctly more complicated than the discrete-time setup with independent and identically distributed data (see Fermanian and Scaillet (2002) for related work in the time series setting). Fourth, we deliver an implementable inference procedure by designing a Newey and West (1994)-type heteroscedasticity and autocorrelation consistent estimator of the long-run asymptotic variance of our realized copula of volatility.
@@ 2 The setting
where $0<x,y<\infty$, while $V_{t}^{X}=(\sigma_{t}^{X})^{2}$ and $V_{t}^{Y}=(\sigma_{t}^{Y})^{2}$ are point-in-time variances of $X$ and $Y$, and $\mathbbm{1}_{\left\{\cdot\right\}}$ is the indicator function. $H_{T}(x,y)$ measures the relative amount of time in $[0,T]$ that the stochastic volatility processes spent at different levels over their support. Hence, (2.2) is the pathwise counterpart of the cumulative distribution function of volatility. Without the $T$ normalization, $H_{T}(x,y)$ is a multivariate extension of the volatility occupation measure studied in Li et al. (2013), which corresponds to the marginal empirical distribution function $F_{T}(x)=H_{T}(x,\infty)$ and $G_{T}(y)=H_{T}(\infty,y)$.
@@ 3 High-frequency estimation
In practice, $V^{X}$ and $V^{Y}$ are not directly observed, so neither is the empirical distribution function of volatility nor the copula transformation. Our empirical strategy is based on the premise that we can approximate the spot variance processes with a realized measure constructed from discretely observed high-frequency data of $X$ and $Y$ over small time intervals. In the asymptotic theory, we are going to both increase the number of blocks and shrink the timespan of each block, while padding it with an increasing number of observations, such that we attain a better and better assessment of the volatility’s path. In doing so, we control for the price jump component. We also deal with the drift, but this is easier since it is asymptotically negligible in our setting. [^8] Christensen et al. (2022) propose a drift burst model, in which the drift term is locally of larger order than the volatility component.
We assume that $(X,Y)$ are sampled at equidistant time points $(i\Delta_{n})_{i=0}^{n}$, where $\Delta_{n}$ is the time gap and $n=\lfloor T/\Delta_{n}\rfloor$ denotes the number of observations over $[0,T]$. We define an increment of the process $Z$ between $(i-1)\Delta_{n}$ and $i\Delta_{n}$ as $\Delta_{i}^{n}Z=Z_{i\Delta_{n}}-Z_{(i-1)\Delta_{n}}$, for $i=1,\dots,n$. Then, as in Li et al. (2013), we adopt the following estimator of $V_{t}^{Z}$ (at time $t$):
@@
The univariate version can be retrieved as $\widehat{F}_{n,T}(x)=\widehat{H}_{n,T}(x,\infty)=\frac{1}{T}\int_{0}^{T}\mathbbm{1}_{\left\{\hat{V}_{t}^{X}\leq x\right\}}\mathrm{d}t$ and $\widehat{G}_{n,T}(y)=\widehat{H}_{n,T}(\infty,y)=\frac{1}{T}\int_{0}^{T}\mathbbm{1}_{\left\{\hat{V}_{t}^{Y}\leq y\right\}}\mathrm{d}t$. The latter are again equivalent to those proposed in Li et al. (2013).
@@ 4 Asymptotic theory
Assumption 3 states that the volatility is an Itô semimartingale. In contrast to the log-price process, this is not a consequence of the no-arbitrage principle. However, it is commonly assumed, because it restricts the local behavior of volatility that is necessary to apply standard estimates for semimartingales in the high-frequency paradigm (see Jacod and Protter, 2012). For example, if $\sigma^{Z}$ is a diffusion, the assumption follows by an application of Itô’s Lemma under certain smoothness conditions. Either way, it is fulfilled for most stochastic volatility models applied in practice, even when volatility has sources of randomness not captured by $Z$. The ability to correlate the Brownian motions of $Z$ and $\sigma^{Z}$ is particularly relevant for financial applications that involve equity data, in order to capture the so-called leverage effect (Black, 1976; Christie, 1982). Note that (4.4) does rule out volatility models driven by a fractional Brownian motion, which feature prominently in some strands of research (e.g., Comte and Renault, 1998; Gatheral et al., 2018). The main difficulty is that in the “rough” regime, the volatility behaves very erratically over short time intervals, and this makes it harder to control the discretization error (see Chong and Todorov, 2023). Christensen et al. (2019) allow $\sigma^{Z}$ to satisfy a Hölder-type condition, in the expectation sense. It may be possible to extend our theoretical framework in this direction as well, but we leave it for a future endeavor.
@@
Assumption 4 is essential in deriving the asymptotic distribution theory. It corresponds to Assumption 5 in Christensen et al. (2019), but for the two-dimensional case, and is again a weak regularity condition that can be verified for most stochastic volatility models. The stationary condition makes it meaningful to speak of a time-invariant distribution function of volatility, $H(x,y)=\mathbb{P}(V_{0}^{X}\leq x,V_{0}^{Y}\leq y)$, from which $F(x)=\mathbb{P}(V_{0}^{X}\leq x)$ and $G(y)=\mathbb{P}(V_{0}^{Y}\leq y)$ can be extracted. Sklar’s theorem then guarantees existence, and uniqueness under further conditions, of the copula $C(u,v)=\mathbb{P}(F(V_{0}^{X})\leq u,G(V_{0}^{Y})\leq v)$. The question is then to what extent $\widehat{H}_{n,T}(x,y)$, consistently estimating $H_{T}(x,y)$ as $\Delta_{n}\rightarrow\infty$, also provides a good proxy of the stationary distribution $H(x,y)$, since there is only a single realization to work with. This, at a bare minimum, additionally requires $T\rightarrow\infty$, so sufficient conditions to apply the ergodic theorem on the volatility path are needed to prompt the law of large numbers for consistency. This is implied by the strong mixing condition, and the polynomial decay of $\alpha_{t}$ ensures that volatility persistence decays suitably fast to derive the limiting distribution of the statistic. In this regard, we should note that since our CLT is derived for a sequence of bounded random variables, all moments exist, so the usual moment condition “trade-off” in Davydov’s inequality is irrelevant, and hence the requirement for the mixing coefficient function is the weakest possible and the autocorrelation function merely needs to be integrable.
@@
This follows Assumption 1 in Christensen et al. (2019) and Andersen et al. (2019). It imposes a light-tailedness condition on the distribution of the various coefficient processes driving the continuous part of $X$ and $Y$, and the large price jumps, but not for the jumps of volatility. It is a common assumption in the long-span setting. The existence of moments of any order is more than we require, but it streamlines the exposition. The technical reason for this condition is that the standard estimates for the increments of semimartingales based on the localization procedure of Jacod and Protter (2012, Section 4.4.1) presumes $T$ is fixed. The above assumption facilitates the derivation of such estimates over $0,\infty)$.
@@
This combines Assumption 2 of Christensen et al. (2019) with Assumption B of Li et al. (2013). The smoothness condition is used to derive a uniform upper bound for estimation error of the empirical copula function. The requirement $\partial_{\bullet}H_{T}(x,y)\neq 0$ rules out “flat spots” in the marginal distribution to ensure that $H_{T}(x,y)\mapsto{C}_{T}(u,v)$ is smooth and non-singular, so that we can apply the functional delta method without getting a degenerate limit process with zero variance. We note that the boundedness in expectation is weaker than asking $\mathbb{E}[\sup_{(x,y)\in\mathbb{R}_{+}^{2}}\partial_{\bullet}H_{T}(x,y)]\leq C$, since the latter requires almost sure control of the pathwise behavior of volatility.
@@
This is identical to the conditions in Theorem 4 of Li et al. (2013), except for the appearance of $\sqrt{T}$ since we operate in the long-span setting. It implies that the volatility error, $\hat{V}_{t}^{Z}-V_{t}^{Z}$, is dominated by $T$, so the high-frequency discretization part is asymptotically negligible. The four terms in $d_{n}$ and $d_{n}^{\prime}$ represent the various sources of the estimation error in $H_{n,T}(x,y)$. Specifically, the first term captures the order of the truncation procedure, induced by replacing truncated increments by those of the continuous part. It depends on whether the jump component has finite or infinite variation. The second and third terms capture the sampling variability and the discretization bias in approximating the spot volatilities. The last order is for the collection of summands over the sampling intervals that contain jumps (with a size larger than some level) in the volatility. The choices of $\varpi,r$, and $\gamma$ imply $d_{n}\rightarrow 0$ and $d_{n}^{\prime}\rightarrow 0$. Thus, we can assume $T\rightarrow\infty$ such that Assumption 7 holds. In particular, for $r\in(0,\frac{1}{2})$ and $\tilde{r}=0$, we can take $\varpi\in(\frac{3}{8-4r},\frac{1}{2})$ and $\gamma=1/2$, from which it follows that $d_{n}=O(\Delta_{n}^{\frac{1}{4}-\iota})$. In this case, $T=O(\Delta_{n}^{-1/4+\epsilon})$ with $\epsilon>0$ verifies the requirement.
The next theorem delivers the functional CLT of the estimation procedure for the stationary distribution function of volatility and the associated copula. It is grounded in a double asymptotic setting with $\Delta_{n}\rightarrow 0$ and $T\rightarrow\infty$ and extends Theorem 1 in Li et al. (2013) and Theorem 3.1 in Christensen et al. (2019).

### 21 Apr 2026 v2

@@ Abstract
We study a new measure of codependency in the second moment of a continuous-time multivariate asset price process, which we name the realized copula of volatility. The statistic is based on local volatility estimates constructed from high-frequency asset returns and affords a nonparametric estimator of the empirical copula of the latent stochastic volatility. We show consistency of our estimator with in-fill asymptotic theory, either with a fixed or increasing time span. In the latter setting, we derive a functional central limit theorem for the empirical process associated with the measurement error of the time-invariant marginal copula of volatility. We also develop a goodness-of-fit test to evaluate hypotheses about the shape of the latter. In a simulation study, we demonstrate that our estimator is a good proxy of both the empirical and marginal copula of volatility, even with a moderate amount of high-frequency data recorded over a relatively short sample. The goodness-of-fit test is found to exhibit size control and excellent power. We implement our framework on high-frequency transaction data from futures contracts that track the U.S. equity and treasury bond market. A Gumbel copula is found to offer a near-perfect bind between the realized variance processes in these data.
@@ 1 Introduction
Modeling the comovement of stochastic volatility across several assets is paramount to risk measurement and management, portfolio allocation, and multi-asset and derivative pricing. It has long been recognized that financial markets are interconnected, causing asset return volatilities—and correlations—to surge in tandem during financial distress, triggering significant volatility spillover effects, posing a potential systemic risk (see, e.g., Diebold and Yilmaz, 2009; Billio, Getmansky, Lo, and Pelizzon, 2012, and references therein). An understanding of volatility codependency is also important for pricing of volatility risk (i.e., the variance risk premium) across asset classes (Carr and Wu, 2009; Bollerslev and Todorov, 2011). The importance of the second moment structure of asset returns is further fueled by the so-called leverage effect that captures a negative association between return and volatility (Black, 1976; Christie, 1982). Thus, from a risk aggregation view the relevant target is often the dependence structure in second moment of returns (namely volatilities and correlations), rather than in the raw returns.
@@
In practice, the assessment of volatility comovements is complicated by the fact that volatility is unobserved, so it requires an empirical route based on an observable proxy. It therefore remains a relatively unexplored area. In recent years, however, the availability of high-frequency data has made it possible to estimate volatility nonparametrically with the realized variance (e.g., Andersen and Bollerslev, 1998; Barndorff-Nielsen and Shephard, 2002). Most of this literature concentrates on estimation of integrated measures of variation, defined as an integral of a smooth function of the latent spot variance process (see Jacod and Rosenbaum, 2013). However, as Li, Todorov, and Tauchen (2013) point out, the mapping from the probability distribution of spot volatility to the distribution of integrated volatility is not one-to-one. Thus, the “integrated” approach may be partially informative about some distributional aspects of volatility and codependency, but it faces a limitation due to the inherent “identification failure.” This may weaken, or outright eliminate, power for testing against relevant alternatives. They suggest to work directly with the pathwise distribution of spot volatility via a so-called occupation measure (e.g., Geman and Horowitz, 1980), which recovers the empirical distribution function of the point-in-time volatility process (see also Li, Todorov, and Tauchen, 2016a, b). [^1] In addition, Christensen, Thyrsgaard, and Veliyev (2019) and Zhang, Li, and Bollerslev (2022) made the univariate framework robust to microstructure noise. Regardless, the volatility occupation literature is entirely univariate and has concentrated on the marginal distribution, rather than looking at the interrelationship between various volatility processes.
@@
The starting point of our analysis is the remarkable result from probability theory known as Sklar’s theorem (Sklar, 1959). It states that for any continuous random vector $(X_{1},\dots,X_{d})$ with cumulative distribution function $H(x_{1},\dots,x_{d})=\mathbb{P}(X_{1}\leq x_{1},\dots,X_{d}\leq x_{d})$, there exists a unique function $C:[0,1]^{d}\rightarrow[0,1]$ called a copula, such that $H(x_{1},\dots,x_{d})=C(H_{1}(x_{1}),\dots,H_{d}(x_{d}))$, where $H_{i}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ is the marginal distribution of $X_{i}$. [^2] The concise mathematical definition is as follows: A $d$-dimensional copula $C:[0,1]^{d}\rightarrow[0,1]$ is the cumulative distribution function of a random vector $(U_{1},\dots,U_{d})$, where the marginal distribution of $U_{i}$ is uniform, i.e. $C(u_{1},\dots,u_{d})=\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ with $\mathbb{P}(U_{i}\leq u_{i})=u_{i}$, for $i=1,\dots,d$ and $0\leq u_{i}\leq 1$. Conversely, given $C(u_{1},\dots,u_{d})$ and $H_{1}(x_{1}),\dots,H_{d}(x_{d})$, $H(x_{1},\dots,x_{d})=C(H_{1}(x_{1}),\dots,H_{d}(x_{d}))$ is a probability distribution. Copulas provide a marginal-free representation of dependence in a multivariate setting, and they enable the researcher to model the marginal features of the data separately. [^3] The term copula originates from the Latin word for “bond” or “tie”. Indeed, many measures of concordance (and discordance), such as Kendall’s tau and Spearman’s rho, can be expressed entirely in terms of the copula (see Section 6). The latter are often more informative than the Pearson’s coefficient of correlation, which captures linear dependence. As emphasized by Embrects, McNeil, and Straumann (1999), correlation is insufficient for risk management. This is explained by the fact that financial markets typically exhibit nonlinear dependence and, further, may possess a high degree of tail concentration (or corner clustering). [^4] The application of copula theory had a revival in financial economics after Li (2000) proposed to model the default correlation in portfolios of bond loans with a Gaussian copula. This enabled tractable analysis of structured credit products (such as collateralized debt obligations). The model became a standard pricing tool on Wall Street. However, since the Gaussian copula has no tail dependence, it tends to understate the probability of concurrent defaults, as seen during the financial crisis (see, for instance, Salmon, 2009, for a critique of the approach). A comprehensive review of the application of copula theory to economic and financial time series is Patton (2012). Therefore, copulas—or descriptive statistics derived from them—are a better representation of dependence. Another advantage of rank-based statistics is their invariance to monotonic transformations (that do not alter the ordering of the data). However, a model-free technique for analyzing the multivariate dependence structure of volatility is still lacking. In particular, tools for estimating and conducting inference on the copula of volatility are not available.
@@
We close that gap and propose to study the dependence structure of several stochastic volatility processes. We make a four-fold contribution to the literature. First, we extend Li, Todorov, and Tauchen (2013) to the multivariate setting. We operate within a general, albeit standard, continuous-time arbitrage-free Itô semimartingale framework for a (multi-dimensional) log-price process, where the random volatility can exhibit rather unrestricted dynamics. We replace the latent spot variance processes by localized, jump-robust, realized variance measures constructed from discretely observed high-frequency returns over short time intervals. The estimation errors are controllable by the sampling interval and bandwidth selection (e.g., Jacod and Protter, 2012). We allow for very active jumps in the price process. The latter can be regarded as “nuisance parameters” in our setting, and we handle them with a truncation device (e.g., Mancini, 2009). Second, we construct a nonparametric estimator of volatility codependency, called the realized copula of volatility, enabling a meaningful cross-asset comparison irrespective of the marginal distribution of the individual volatility processes. Third, we develop the necessary large-sample theory for our estimator aligned with its target at each horizon. In the in-fill asymptotic limit with fixed time span, we show the realized copula of volatility converges, uniformly, in probability to its empirical counterpart that captures the volatility dependency over the time interval observed so far. Importantly, we allow for nonstationary volatility. In a double asymptotic in-fill and long-span setting—with stationary and weakly dependent volatility—we establish a functional central limit theorem in the form of weak convergence of the measurement error of the realized volatility copula. We show how this can be exploited to construct pointwise or, with a bit more effort, uniform confidence bands. A goodness-of-fit test for assessing the shape of the volatility copula is also proposed. The latter procedure can, for instance, be used to gauge the adequacy of various parametric copulas; an approach we look further at in our application. This builds upon and extends earlier results on the empirical copula process (Deheuvels, 1979; Fermanian, Radulović, and Wegkamp, 2004). However, as explained above, our framework is distinctly more complicated than the discrete-time setup with independent and identically distributed data (see Fermanian and Scaillet (2002) for related work in the time series setting). Fourth, we deliver an implementable inference procedure by designing a Newey and West (1994)-type heteroscedasticity and autocorrelation consistent estimator of the long-run asymptotic variance of our realized copula of volatility.
@@ 2 The setting
where $0<x,y<\infty$, while $V_{t}^{X}=(\sigma_{t}^{X})^{2}$ and $V_{t}^{Y}=(\sigma_{t}^{Y})^{2}$ are point-in-time variances of $X$ and $Y$, and $\mathbbm{1}_{\left\{\cdot\right\}}$ is the indicator function. $H_{T}(x,y)$ measures the relative amount of time in $[0,T]$ that the stochastic volatility processes spent at different levels over their support. Hence, (2.2) is the pathwise counterpart of the cumulative distribution function of volatility. Without the $T$ normalization, $H_{T}(x,y)$ is a multivariate extension of the volatility occupation measure studied in Li, Todorov, and Tauchen (2013), which corresponds to the marginal empirical distribution function $F_{T}(x)=H_{T}(x,\infty)$ and $G_{T}(y)=H_{T}(\infty,y)$.
@@ 3 High-frequency estimation
In practice, $V^{X}$ and $V^{Y}$ are not directly observed, so neither is the empirical distribution function of volatility nor the copula transformation. Our empirical strategy is based on the premise that we can approximate the spot variance processes with a realized measure constructed from discretely observed high-frequency data of $X$ and $Y$ over small time intervals. In the asymptotic theory, we are going to both increase the number of blocks and shrink the timespan of each block, while padding it with an increasing number of observations, such that we attain a better and better assessment of the volatility’s path. In doing so, we control for the price jump component. We also deal with the drift, but this is easier since it is asymptotically negligible in our setting. [^8] Christensen, Oomen, and Renò (2022) propose a drift burst model, in which the drift term is locally of larger order than the volatility component.
We assume that $(X,Y)$ are sampled at equidistant time points $(i\Delta_{n})_{i=0}^{n}$, where $\Delta_{n}$ is the time gap and $n=\lfloor T/\Delta_{n}\rfloor$ denotes the number of observations over $[0,T]$. We define an increment of the process $Z$ between $(i-1)\Delta_{n}$ and $i\Delta_{n}$ as $\Delta_{i}^{n}Z=Z_{i\Delta_{n}}-Z_{(i-1)\Delta_{n}}$, for $i=1,\dots,n$. Then, as in Li, Todorov, and Tauchen (2013), we adopt the following estimator of $V_{t}^{Z}$ (at time $t$):
@@
The univariate version can be retrieved as $\widehat{F}_{n,T}(x)=\widehat{H}_{n,T}(x,\infty)=\frac{1}{T}\int_{0}^{T}\mathbbm{1}_{\left\{\hat{V}_{t}^{X}\leq x\right\}}\mathrm{d}t$ and $\widehat{G}_{n,T}(y)=\widehat{H}_{n,T}(\infty,y)=\frac{1}{T}\int_{0}^{T}\mathbbm{1}_{\left\{\hat{V}_{t}^{Y}\leq y\right\}}\mathrm{d}t$. The latter are again equivalent to those proposed in Li, Todorov, and Tauchen (2013).
@@ 4 Asymptotic theory
Assumption 3 states that the volatility is an Itô semimartingale. In contrast to the log-price process, this is not a consequence of the no-arbitrage principle. However, it is commonly assumed, because it restricts the local behavior of volatility that is necessary to apply standard estimates for semimartingales in the high-frequency paradigm (see Jacod and Protter, 2012). For example, if $\sigma^{Z}$ is a diffusion, the assumption follows by an application of Itô’s Lemma under certain smoothness conditions. Either way, it is fulfilled for most stochastic volatility models applied in practice, even when volatility has sources of randomness not captured by $Z$. The ability to correlate the Brownian motions of $Z$ and $\sigma^{Z}$ is particularly relevant for financial applications that involve equity data, in order to capture the so-called leverage effect (Black, 1976; Christie, 1982). Note that (4.4) does rule out volatility models driven by a fractional Brownian motion, which feature prominently in some strands of research (e.g., Comte and Renault, 1998; Gatheral, Jaisson, and Rosenbaum, 2018). The main difficulty is that in the “rough” regime, the volatility behaves very erratically over short time intervals, and this makes it harder to control the discretization error (see Chong and Todorov, 2023). Christensen, Thyrsgaard, and Veliyev (2019) allow $\sigma^{Z}$ to satisfy a Hölder-type condition, in the expectation sense. It may be possible to extend our theoretical framework in this direction as well, but we leave it for a future endeavor.
@@
Assumption 4 is essential in deriving the asymptotic distribution theory. It corresponds to Assumption 5 in Christensen, Thyrsgaard, and Veliyev (2019), but for the two-dimensional case, and is again a weak regularity condition that can be verified for most stochastic volatility models. The stationary condition makes it meaningful to speak of a time-invariant distribution function of volatility, $H(x,y)=\mathbb{P}(V_{0}^{X}\leq x,V_{0}^{Y}\leq y)$, from which $F(x)=\mathbb{P}(V_{0}^{X}\leq x)$ and $G(y)=\mathbb{P}(V_{0}^{Y}\leq y)$ can be extracted. Sklar’s theorem then guarantees existence, and uniqueness under further conditions, of the copula $C(u,v)=\mathbb{P}(F(V_{0}^{X})\leq u,G(V_{0}^{Y})\leq v)$. The question is then to what extent $\widehat{H}_{n,T}(x,y)$, consistently estimating $H_{T}(x,y)$ as $\Delta_{n}\rightarrow 0$, also provides a good proxy of the stationary distribution $H(x,y)$, since there is only a single realization to work with. This, at a bare minimum, additionally requires $T\rightarrow\infty$, so sufficient conditions to apply the ergodic theorem on the volatility path are needed to prompt the law of large numbers for consistency. This is implied by the strong mixing condition, and the polynomial decay of $\alpha_{t}$ ensures that volatility persistence decays suitably fast to derive the limiting distribution of the statistic. In this regard, we should note that since our CLT is derived for a sequence of bounded random variables, all moments exist, so the usual moment condition “trade-off” in Davydov’s inequality is irrelevant, and hence the requirement for the mixing coefficient function is the weakest possible and the autocorrelation function merely needs to be integrable.
@@
This follows Assumption 1 in Christensen, Thyrsgaard, and Veliyev (2019) and Andersen, Thyrsgaard, and Todorov (2019). It imposes a light-tailedness condition on the distribution of the various coefficient processes driving the continuous part of $X$ and $Y$, and the large price jumps, but not for the jumps of volatility. It is a common assumption in the long-span setting. The existence of moments of any order is more than we require, but it streamlines the exposition. The technical reason for this condition is that the standard estimates for the increments of semimartingales based on the localization procedure of Jacod and Protter (2012, Section 4.4.1) presumes $T$ is fixed. The above assumption facilitates the derivation of such estimates over $0,\infty)$.
@@
This combines Assumption 2 of Christensen, Thyrsgaard, and Veliyev (2019) with Assumption B of Li, Todorov, and Tauchen (2013). The smoothness condition is used to derive a uniform upper bound for estimation error of the empirical copula function. The requirement $\partial_{\bullet}H_{T}(x,y)\neq 0$ rules out “flat spots” in the marginal distribution to ensure that $H_{T}(x,y)\mapsto{C}_{T}(u,v)$ is smooth and non-singular, so that we can apply the functional delta method without getting a degenerate limit process with zero variance. We note that the boundedness in expectation is weaker than asking $\mathbb{E}[\sup_{(x,y)\in\mathbb{R}_{+}^{2}}\partial_{\bullet}H_{T}(x,y)]\leq C$, since the latter requires almost sure control of the pathwise behavior of volatility.
@@
This is identical to the conditions in Theorem 4 of Li, Todorov, and Tauchen (2013), except for the appearance of $\sqrt{T}$ since we operate in the long-span setting. It implies that the volatility error, $\hat{V}_{t}^{Z}-V_{t}^{Z}$, is dominated by $T$, so the high-frequency discretization part is asymptotically negligible. The four terms in $d_{n}$ and $d_{n}^{\prime}$ represent the various sources of the estimation error in $H_{n,T}(x,y)$. Specifically, the first term captures the order of the truncation procedure, induced by replacing truncated increments by those of the continuous part. It depends on whether the jump component has finite or infinite variation. The second and third terms capture the sampling variability and the discretization bias in approximating the spot volatilities. The last order is for the collection of summands over the sampling intervals that contain jumps (with a size larger than some level) in the volatility. The choices of $\varpi,r$, and $\gamma$ imply $d_{n}\rightarrow 0$ and $d_{n}^{\prime}\rightarrow 0$. Thus, we can assume $T\rightarrow\infty$ such that Assumption 7 holds. In particular, for $r\in(0,\frac{1}{2})$ and $\tilde{r}=0$, we can take $\varpi\in(\frac{3}{8-4r},\frac{1}{2})$ and $\gamma=1/2$, from which it follows that $d_{n}=O(\Delta_{n}^{\frac{1}{4}-\iota})$. In this case, $T=O(\Delta_{n}^{-1/4+\epsilon})$ with $\epsilon>0$ verifies the requirement.
The next theorem delivers the functional CLT of the estimation procedure for the stationary distribution function of volatility and the associated copula. It is grounded in a double asymptotic setting with $\Delta_{n}\rightarrow 0$ and $T\rightarrow\infty$ and extends Theorem 1 in Li, Todorov, and Tauchen (2013) and Theorem 3.1 in Christensen, Thyrsgaard, and Veliyev (2019).